OpenStudy (anonymous):

Verify the identity. Justify each step.tan theta + cot theta= 1/ sin theta cos theta

OpenStudy (loser66):

what does "justify each step" mean?

OpenStudy (anonymous):

like show how it is correct, i am guessing. @loser66

OpenStudy (anonymous):

\[\frac{sin\theta}{cos\theta}+\frac{cos\theta}{sin\theta}= make~them~have~the~same~denominator~then,~you~\\got~the~answer\]

OpenStudy (anonymous):

try it, please.

OpenStudy (jdoe0001):

$$\bf tan(\theta) + cot(\theta) = \frac{1}{sin(\theta)cos(\theta)}\\ \frac{sin(\theta)}{cos(\theta)}+\frac{cos(\theta)}{sin(\theta)}= \frac{1}{sin(\theta)cos(\theta)}\\ \frac{sin^2(\theta)+cos^2(\theta)}{cos(\theta)sin(\theta)}= \frac{1}{sin(\theta)cos(\theta)} $$